The hail grains during the fall encounter the force of resistance in the air. Once the uniform motion is established, the 2 g hailstone travels 120 m for 20 s. Calculate the coefficient of resistance for the friction force without taking into account the Archimedes force.
1
Expert's answer
2020-12-15T11:47:42-0500

Find the speed at which the force of gravity equals drag force:


"v=\frac dt."

The condition of falling at a constant speed:


"mg=kv,\\\space\\
k=\frac{mg}{v}=\frac{mgt}{d}=1.64\cdot10^{-4}\text{ kg/s}."